OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5).

OpenStudy (anonymous):

(f0g)(-5)=f(g(-5)) g(-5)=-4(-5)+3=23 f(g(-5))=f(20)=-2(20)-7=-40-7=-47

OpenStudy (anonymous):

sorry f(23)=-2(23)-7=-46-7=-53

OpenStudy (anonymous):

thank youuu!!